Rewrite the expression without using the absolute value symbol
|x-2| if x<2

Found 2 solutions by lwsshak3, Theo:
Answer by lwsshak3(11628)

Rewrite the expression without using the absolute value symbol
|x-2| if x<2
The rule is that if the expression within the absolute bars is positive, just remove the bars. Conversely, if negative, put a minus sign in front of the expression after removing the bars.
In the given case, since x<2, (x-2) is negative,therefore, |x-2|= -(x-2)=2-x

Answer by Theo(13342)

In general:

|x-2| = (x-2) if (x-2) is greater than or equal to 0.
|x-2| = -(x-2) if (x-2) is less than 0.

If x is less than 2, then the expression |x-2| will be less than 0, so the expression |x-2| will be equivalent to -(x-2).

Example:

let x = 1.
|1-2| = |-1| = -(-1) = 1